### v1

@@ 4.3 Application Domains
5 Solutions to Engineer Software for QCaaS (RQ-2)
@@
    Service lifecycle: Four activities namely conception, modelling, assembly, and deployment are collectively referred to as foundational activities for IBM SOA lifecycle [26]. These four generic activities are used to contextualise and organise seven steps to engineer software services for quantum computing. Specifically, an activity guides about what needs to be done? whereas, the step illustrates how it is to be done? For example in study [S1], the first activity namely conception provides foundations for system analysts to conceive the idea of developing a service to execute mathematical computations. Functional aspects and quality attributes as two specific steps of conception can guide service developers to consider the required functionality (e.g., quantum service delivery) and quality (e.g., dynamic adaptation) that needs to ensured in the implemented service. Each of the seven steps to engineer QCaaS are organised under four activities of SOA lifecycle and presented in Table 2.
@@
    Extracted data from studies: The data extracted from primary studies recorded in represents the published evidence that is synthesised and summarised in Table 2. Table 2 acts as a structured catalogue, detailing each reviewed study that represents published evidence (Study ID) and seven steps of QCaaS under four service lifecycle activities. For example, Table 2 can be looked up to quickly identify that study referred to as [S1] focuses on (i) conception as optimal delivery of quantum service (functional aspect) via a runtime dynamic adaptation and selection of a specific service (quality attribute) from a pool of available services. To support (ii) modelling of this service UML deployment diagram is used as a notation and API Gateway pattern is applied to manage client request and runtime selection of most suitable service. To (iii) assemble this service, a usecase on process automation and service optimisation is selected and Flask as a Python-based framework is used to implement the service. Finally, for (iv) deployment of the service Amazon Braket is used as a service delivery platform.
@@
    Illustrative example: We also exemplify Shor’s algorithms as an example to architect quantum service-orientation. The example provides a scenario to and a step-by-step illustration and elaboration of the four phases and seven activities of service lifecycle to engineer QCaaS, presented in Figure 8.
@@ 5.1.1 Functional Aspects
Based on our review, we identified and classified functional aspects into five (5) specific types while quality attributes are classified into seven (7) specific types as highlighted in Figure 5 a). These five types include Continuous Deployment (identified in a total of 3 out of 41 (3/41) reviewed studies that represent 7% approx. of reviewed literature), Quantum Experimental Services (5/41, 12.5%), Quantum Cloud and Service Orchestration (5/41, 12.5%), Quantum-Classic Hybrid Computing (4/41, 10%), Quantum Service Delivery (2/41, 51%), and Miscelleneous (3/41, 7%). In the context of quantum service patterns, some stuides such as [S10][S16] refer to the qnatum-classic hybrid comoputing as qunatum-classic split pattern. For the sake of clarity, the quantum-classic hybrid computing is a general reference to the computing model, whereas quantum-classic split is a specific pattern to implement the hybrid computing model. In the following, we briefly introduce each specific type of functional aspects and quality attributes and exemplify one of them as a representative example for elaboration. For example, considering ‘quantum cloud and service orchestration’ type, the studies [S32][S38] focus on orchestrating software services that can manipulate QC hardware that is deployed on a remote server and made available via the cloud computing model. Specifically, [S38] allows QC vendors to offer quantum resources (e.g., QuBits) that can be utilised via on-demand plans and pay-per-use model, as illustrated in Figure 2. As part of service conception, we summarise the functional aspects as:
  • Continuous deployment enabled via quantum micro servicing is identified in a total of 3 out of 41 studies (representing 7% of all reviewed studies). By following continuous deployment, the functionality of a service is developed, tested, and deployed for a seamless delivery of software services that can be executed on QC platforms.
@@ 5.1.2 Quality Attributes
  • Computational efficiency refers to achieving optimal throughput and optimizing execution time etc. for scalable and high-performance computing as reported in 18 studies. In the context of computational efficiency for QCs, the term quantum computational supremacy is often used, referring to a computation threshold where a quantum computer is computationally efficient and outperforms the optimal classical supercomputers in solving a problem [2] [2]. For example, the study [S18] proposes a quantum orchestrator for load-balancing and efficient execution of quantum tasks on platforms like Amazon Braket and IBM Quantum, supporting both quantum and classical workflows.
  • Service evolution refers to the maintainability and adaptability of software services, ensuring their continuous improvement and integration. A total of seven studies have explored this aspect. To provide a more fine-grained analysis, we classify service evolution into two main types: (1) Dynamic adaptation (i.e., adaptive evolution), which enables services to adjust to changing requirements, and (2) Service integration and interoperability (i.e., perfective evolution), which focuses on enhancing compatibility and seamless interaction between services. In the context of quantum software services, service evolution is particularly critical due to the rapid advancements in quantum algorithms, hardware, and hybrid classical-quantum systems [S7]. Adaptive evolution ensures that quantum services remain efficient despite evolving quantum computing paradigms, while perfective evolution facilitates the integration of quantum services with existing classical infrastructures, promoting broader adoption and interoperability [S3].
@@
  • Security in the context of quantum service-orientation refers to services that enable secure management and execution of systems and software executed on QC platforms - often referred to as security as a service [S14][S23]. Data security solutions offered via quantum services include robust encryption, quantum-safe cryptography, and resilience against quantum attacks, safeguarding data integrity, confidentiality, and secure communication in quantum-classical hybrid computing environments. Security as a critical quality attribute is identified in a total of 4 studies [S9][S11][S14][S23]. For example, the study [S23] proposes a smart city security architecture that can be deployed on various QC platforms, such as D-Wave and IBM Quantum Platform for real-time data classification and addressing security risks for smart city devices.
@@ 5.2 Activity II - Service Modelling
After conception, the modelling activity relies on notations as a means to represent the design and patterns as best practices to create service design based on functional aspects and quality attributes. Modeling notations include but are not limited to graphical, descriptive, or mathematical representation of service design. Patterns complement the notations, as artifacts of reuse knowledge, to model an optimal design. As a specific example, study [S3] uses UML class and sequence diagrams as UML-based modelling notations to represent quantum search algorithms. The quantum-classic split as a layered pattern is applied to modularise the search algorithm in terms of implementing the orchestration of classic and quantum computations.
@@ 5.2.1 Modelling Notations
These are fundamental to the creation, maintenance, and evolution of models such as ontological structures and graph-based diagrams that provide a visual representation, whereas architectural description languages support a textual specification for software-intensive systems [52] [53]. Recent trends in software engineering that promote model-driven and low-code application development have resulted in transitioning developers’ focus from coding to software modeling for implementation [54]. Low-code application development process leverages the principle and practices of model-driven engineering to utilise model(s) as first-class entities in software development [54]. Moreover, with a recent emergence of large language models as software development bots, there is an increasing reliance on high-level descriptive models or graphical representation to generate source code for qunatum algorithms [55]. Investigating software models and modeling notations that help create service models is essential to support model-driven perspective to QSE, consequently facilitating quantum code developers to abstract implementation-specific complexities, via model-driven QSE, while developing quantum services [52][54]. For example, the study [S3] reports a class diagram as a structural view of the system to model the attributes and methods of entities (user, service provider, authentication etc.) of quantum computing as a service. This SMS indicates five main types of notations to model services in QCaaS detailed below based on the presentation in Figure 5.
  • Process models in service development lifecycle are used to visualise the workflows, activities, and interactions to analyze, design, and optimise services, ensuring an alignment with business goals, user needs, and requirements implemented as software services. Process models as a modeling notation are identified in a total of 3 studies [S2][S16][S17] that use a process-centric view and business process modelling notation to create service design. For example, in order to abstract out complex and implementation-specific details of quantum algorithmic implemenations, the study [S17] aims to facilitate developers to concentrate on application development rather than low-level quantum hardware details. In order to achieve this, the authors utilise the business process modeling notation, the well-known BPMN to design a process model that is used for generating quantum service source code from API specifications and quantum circuits.
@@
  • Unified Modelling Language (UML) based models are represented via a multitude of notations, such as class and component diagrams that represent the structure [S3], while sequence and deployment diagrams [S6][S13] are used to represent the runtime or behavioural view of software services. UML diagrams and profiles represent the status-quo in software modeling and are seen as the de-facto notation in the software and service developers’ community to model classical software with growing adoption in QSE [52]. For example, the study [S10] proposes a reference architecture for quantum computing as a service. A proof-of-the-concept design and implementation of qunatum software services is provided via the Shor’s algorithm that is modeled using UML sequence diagram (modeling algorithmic behaviour) and UML deployment diagram (showing the qunatum-classic split) based deployment of the service. The role of UML-based modeling is further exemplified in the context of pattern-based modeling of quantum software services - illustrated in Figure 6.
  • Workflow model for service design represents the task to be performed (e.g., functional aspects), roles (entities or actors that interact with the tasks), and interactions to guide the structured development of software. In the reviewed studies, often the terms process modeling and workflow modeling have been used interchangeably, generally referring to modeling a flow of service execution [S16][S34]. From a technical perspective, a distinction between the two needs to be maintained. Specifically, a business process model provides a high-level view of organisational activities, focusing on goals, roles, and processes that can be automated via several services. In contrast, a workflow model details the specific sequence of tasks and data flows, often expressed as flow charts, emphasising service execution and automation. For example, the study [S34] uses workflow modeling to represent the execution and interaction of services to design, implement, and deeply qunatum algorithms.
@@ 5.2.2 Design Patterns
Patterns [^2] In software design and architecture context, the terms pattern and style are often used interchangeably to refer to best practices or frequently used solutions for recurring software modelling challenges. However, for technical clarity, a style refers to a broad structural blueprint that defines a family of systems by their components and relationships (e.g., client-server style), whereas a pattern is a reusable solution to a specific architectural problem, often addressing component interaction or deployment within a given style (e.g., a layered pattern used to implement the client-server style). represent a concentrated wisdom of software designers as reusable knowledge and best practices that can be leveraged to address recurring problems of software design and architecture. Considering a lack of professional expertise in QSE (e.g., quantum domain engineers, quantum algorithm designers, quantum software architects, etc.) patterns as artifacts of reuse can help novice developers during quantum software development to rely on existing best practices [56]. This SMS highlights that the literature on QCaaS reports five patterns, namely the API Gateway, Layered Architecture, Classic-Quantum Split, Service Façade, and Repository Pattern as shown in Figure 5. Figure 6 depicts pattern thumbnails [57] that provide a concise visual summary of the pattern, typically including its name and structural representation of pattern elements and their interactions.
@@
Example: API Gateway for Amazon Braket enables clients to access QC resources via a Gateway - a pattern used by Amazon as a platform agnostic solution to intermediate between QC users and QC vendors via Amazon Braket [58]. For example, in Figure 6 a), the thumbnail suggests a pattern named Quantum API Gateway as a layered system comprises of three main elements including a) Quantum Service Clients that can invoke available services on b) Quantum Server via routing and recommendation by c) Quantum API Gateway, detailed in [S1]. Figure 6 a) illustrates Amazon Braket that is a managed service provided by Amazon Web Services (AWS), using the API Gateway pattern, for accessing QC resources via QCaaS. The pattern comprises three key layers. At the client node, users aim to access QC resources at the client node, having Jupyter Notebook and AWS Console, enabling them to design, simulate, and submit quantum programs via the Braket SDK. The gateway service, i.e., Amazon Braket, acts as an intermediary that abstracts hardware complexity and manages execution workflows, including device selection, queuing, and error handling. Finally, the quantum computing resources layer connects users to multiple quantum hardware providers (e.g., Vendor A, Vendor B, etc.). This model provides a platform-independent interface, facilitating seamless experimentation across various quantum technologies. By decoupling development from hardware integration, Amazon Braket supports scalable, vendor-neutral quantum computing for researchers, developers, and organisations exploring quantum applications.
@@
Pattern documentation: One of the most common mechanisms to document and store patterns are pattern templates and pattern catalogues that provide structured management of available patterns [57]. Our intent is not to provide comprehensive details on pattern documentation and management, therefore, we rely on pattern thumbnails from Figure 6 to concisely summarise the identified patterns. We briefly introduce each pattern and elaborate on the Quantum Classic Split pattern, with an example, detailed below.
@@
  • Quantum-Classic Split is identified in a total of 3 studies is a quantum version of the Splitter pattern and enables hybrid computation that can be divided between classical and quantum computations [S10][12][s24]. As shown in Figure 6 c), the classic-quantum split is driven by the workflow, also referred to as the split logic to divide the execution of computations. Figure 6 shows that the qunatum-classic workflow modelled as a UML activity computes and delegates activities to be executed on classical and quantum computer. For example, the component and connector architectural view in Figure 6 e) shows that a ClassicalComponent and QuantumDriver deployed on a classical machine can invoke QuantumComponent on a quantum computer. A typical real-world example of the Classic-Quantum Split pattern is Shor’s algorithm for prime factorisation. The algorithm divides tasks between classical and quantum computations such that classical processing handles choosing a random integer, and performing the greatest common divisor checks, and postprocessing, whereas the quantum part efficiently determines the period using modular exponentiation and Quantum Fourier Transform (QFT). This hybrid approach leverages quantum speedup for order finding while relying on classical hardware for the remaining computations.
@@
  • Service Façade for quantum software services refers to a classical interface that abstracts and simplifies access to complex quantum functionalities. It acts as a bridge between classical clients and quantum backends, hiding the complexities of quantum computing — such as circuit design, qubit management, and backend selection — behind a well-organised and easy to implement service-oriented interfaces. Service façade is identified into a total of 2 studies. For example, the study [S7] utilises the service façade pattern, which wraps the classical (legacy software services) under the interfaces of quantum microservices that can be executed on Amazon Braket platform.
@@
L Quantum service modeling supports a model-based representation (i.e., creating design or architecture models) of functionality and quality of the conceived services. Modelling phase is supported via: – Modelling notations can assist software engineers to shift their focus from complex implementation (low-level source-coding) specific details to a design perspective (high-level system view). Modeling can incrementally transform functional aspects to service models, leading to service implementation via model-driven engineering or low-code development. The literature suggests that the modeling phase needs to leverage existing or innovative notations, such as UML diagrams and quantum circuit modells to design quantum software services. – Patterns complementăsoftware models by providing reuse knowledge and best practices of classical and quantum SE during service modelling. The review has identified several classical as well as quantum-specific patterns such as Quantum API Gateway and Qunatum-Classic Split that can facilitate architects to design quantum software services.

### v2

@@ 4.3 Application Domains
5 Solutions to Engineer QCaaS (RQ-2)
@@
    Service lifecycle: Four activities namely conception, modelling, assembly, and deployment are collectively referred to as foundational activities for IBM SOA lifecycle [28]. These four generic activities are used to contextualise and organise seven steps to engineer software services for quantum computing. Specifically, an activity guides about what needs to be done? whereas, the step illustrates how it is to be done? For example in study [S1], the first activity namely conception provides foundations for system analysts to conceive the idea of developing a service to execute mathematical computations. Functional aspects and quality attributes as two specific steps of conception can guide service developers to consider the required functionality (e.g., quantum service delivery) and quality (e.g., dynamic adaptation) that needs to be ensured in the implemented service. Each of the seven steps to engineer QCaaS are organised under four activities of SOA lifecycle as in Table 2.
@@
    Extracted data from studies: The data extracted from primary studies represents the published evidence that is synthesised and summarised in Table 2. Table 2 acts as a structured catalogue, detailing each reviewed study that represents published evidence (Study ID) and seven steps of QCaaS under four service lifecycle activities. For example, Table 2 can be looked up to quickly identify that study referred to as [S1] focuses on (i) conception as optimal delivery of quantum service (functional aspect) via a runtime dynamic adaptation and selection of a specific service (quality attribute) from a pool of available services. To support (ii) modelling of this service UML deployment diagram is used as a notation and API Gateway pattern is applied to manage client request and runtime selection of most suitable service. To (iii) assemble this service, a usecase on process automation and service optimisation is selected and Flask as a Python-based framework is used to implement the service. Finally, for (iv) deployment of the service, Amazon Braket is used as a service delivery platform.
@@
    Illustrative example: We also exemplify Shor’s algorithms as an example to architect quantum service-orientation. The example provides a scenario for a step-by-step illustration and elaboration of the four phases and seven activities of service lifecycle to engineer QCaaS, presented in Figure 8.
@@ 5.1.1 Functional Aspects
Based on our review, we identified and classified functional aspects into five (5) specific types while quality attributes are classified into seven (7) specific types as highlighted in Figure 5 a). These five types include Continuous Deployment (identified in a total of 3 out of 41 (3/41) reviewed studies that represent 7% approx. of reviewed literature), Quantum Experimental Services (5/41, 12.5%), Quantum Cloud and Service Orchestration (5/41, 12.5%), Quantum-Classic Hybrid Computing (4/41, 10%), Quantum Service Delivery (2/41, 51%), and Miscelleneous (3/41, 7%). In the context of quantum service patterns, some stuides such as [S10][S16] refer to the quantum-classic hybrid computing as quantum-classic split pattern. For the sake of clarity, the quantum-classic hybrid computing is a general reference to the computing model, whereas quantum-classic split is a specific pattern to implement the hybrid computing model. In the following, we briefly introduce each specific type of functional aspects and quality attributes and exemplify one of them as a representative example for elaboration. For example, considering ‘quantum cloud and service orchestration’ type, the studies [S32][S38] focus on orchestrating software services that can manipulate QC hardware that is deployed on a remote server and made available via the cloud computing model. Specifically, [S38] allows QC vendors to offer quantum resources (e.g., QuBits) that can be utilised via on-demand plans and pay-per-use model, as illustrated in Figure 2. As part of service conception, we summarise the functional aspects as:
  • Continuous deployment enabled via quantum microservicing and is identified in a total of 3 out of 41 studies (representing 7% of all reviewed studies). By following continuous deployment, the functionality of a service is developed, tested, and deployed for a seamless delivery of software services that can be executed on QC platforms.
@@ 5.1.2 Quality Attributes
  • Computational efficiency refers to achieving optimal throughput and optimising execution time etc. for scalable and high-performance computing, reported in 18 studies. In the context of efficiency for QCs, the term quantum computational supremacy is often used, referring to a computation threshold where a quantum computer is computationally efficient and outperforms the optimal classical supercomputers in solving a problem [2]. For example, the study [S18] proposes a quantum orchestrator for load-balancing and efficient execution of quantum tasks on platforms like Amazon Braket and IBM Quantum, supporting both quantum and classical workflows.
  • Service evolution refers to the maintainability and adaptability of software services, ensuring their continuous improvement and integration. A total of seven studies have explored this aspect. To provide a more fine-grained analysis, we classify service evolution into two main types: (1) Dynamic adaptation (i.e., adaptive evolution), which enables services to adjust to changing requirements, and (2) Service integration and interoperability (i.e., perfective evolution), which focuses on enhancing compatibility and seamless interaction between services. In the context of quantum software services, service evolution is particularly critical due to the rapid advancements in quantum algorithms, hardware, and hybrid quantum-classic systems [S7]. Adaptive evolution ensures that quantum services remain efficient despite evolving quantum computing paradigms, while perfective evolution facilitates the integration of quantum services with existing classical infrastructures, promoting broader adoption and interoperability [S3].
@@
  • Security in the context of quantum service-orientation refers to services that enable secure management and execution of systems and software executed on QC platforms - often referred to as security as a service [S14][S23]. Data security solutions offered via quantum services include robust encryption, quantum-safe cryptography, and resilience against quantum attacks, safeguarding data integrity, confidentiality, and secure communication in quantum-classical hybrid computing environments. Security as a critical quality attribute is identified in a total of 4 studies [S9][S11][S14][S23]. The study [S23] proposes a smart city security architecture that can be deployed on various QC platforms, such as D-Wave and IBM Quantum Platform for real-time data classification and addressing security risks for smart city devices.
@@ 5.2 Activity II - Service Modelling
After conception, the modelling activity relies on notations as a means to represent the design and patterns as best practices to create service design based on functional aspects and quality attributes. Modeling notations include but are not limited to graphical, descriptive, or mathematical representation of service design. Patterns complement the notations, as artifacts of reuse knowledge, to model an optimal design. As a specific example, study [S3] uses UML class and sequence diagrams as UML-based modeling notations to represent quantum search algorithms. The quantum-classic split as a layered pattern is applied to modularise the search algorithm in terms of implementing the orchestration of classic and quantum computations.
@@ 5.2.1 Modelling Notations
These are essential for creating, maintaining, and evolving models such as ontologies and graph-based diagrams, which provide visual representations, while architectural description languages offer textual specifications for software-intensive systems [53] [54]. Recent trends in software engineering, emphasising model-driven and low-code application development, have shifted developers’ focus from traditional coding to software modeling for implementation [55]. The low-code approach applies model-driven engineering principles, treating models as first-class entities throughout the software development process. Moreover, with a recent emergence of large language models as software development bots, there is an increasing reliance on high-level descriptive models or graphical representation to generate source code for qunatum algorithms [56]. Exploring software models and modeling notations that support service modeling is crucial for adopting a model-driven approach in QSE. This approach enables quantum code developers to abstract implementation-specific complexities while creating quantum services [53][55]. For example, study [S3] presents a class diagram as a structural view of the system, modeling attributes and methods of entities such as users, service providers, and authentication in QCaaS. This SMS identifies five main types of notations for modeling services in QCaaS, as illustrated in Figure 5.
  • Process models in service development lifecycle are used to visualise the workflows, activities, and interactions to analyze, design, and optimise services, ensuring an alignment with business goals, user needs, and requirements implemented as software services. Process models as a modeling notation are identified in a total of 3 studies [S2][S16][S17] that use a process-centric view and business process modelling notation to create service design. For example, in order to abstract out complex and implementation-specific details of quantum algorithmic implemenations, the study [S17] aims to facilitate developers to concentrate on application development rather than low-level quantum hardware details. In order to achieve this, the authors of the study utilised the business process modeling notation (BPMN) to design a process model that is used for generating quantum service source code from API specifications and quantum circuits.
@@
  • Unified Modelling Language (UML) based models use a variety of notations inclusing the class and component diagrams that depict structural aspects [S3], while sequence and deployment diagrams capture runtime or behavioral views of software services [S6][S13]. UML diagrams and profiles remain the standard in software modeling and are widely adopted by software and service developers, with increasing application in QSE [53]. For example, the study [S10] proposes a reference architecture for quantum computing as a service. A proof-of-the-concept design and implementation of qunatum software services is provided via the Shor’s algorithm that is modeled using UML sequence diagram (modeling algorithmic behaviour) and UML deployment diagram (showing the qunatum-classic split) based deployment of the service. The role of UML-based modeling is further exemplified in the context of pattern-based modeling of quantum software services - illustrated in Figure 6.
  • Workflow model for service design represents the task to be performed (e.g., functional aspects), roles (entities or actors that interact with the tasks), and interactions to guide the structured development of software. In the reviewed studies, often the terms process modeling and workflow modeling have been used interchangeably, generally referring to modeling a flow of service execution [S16][S34]. From a technical perspective, a distinction between the two needs to be maintained. Specifically, a business process model provides a high-level view of organisational activities, focusing on goals, roles, and processes that can be automated via several services. In contrast, a workflow model details the specific sequence of tasks and data flows, often expressed as flow charts, emphasising service execution and automation. For example, the study [S34] uses workflow modeling to represent the execution and interaction of services to design, implement, and deploy qunatum algorithms.
@@ 5.2.2 Design Patterns
Patterns [^2] In software design and architecture context, the terms pattern and style are often used interchangeably to refer to best practices or frequently used solutions for recurring software modelling challenges. However, for technical clarity, a style refers to a broad structural blueprint that defines a family of systems by their components and relationships (e.g., client-server style), whereas a pattern is a reusable solution to a specific architectural problem, often addressing component interaction or deployment within a given style (e.g., a layered pattern used to implement the client-server style). capture the collective expertise of software designers as reusable knowledge and best practices for addressing recurring design and architectural challenges. Given the limited professional expertise in QSE, such as quantum engineers, algorithm designers, and software architects, these patterns serve as reusable artifacts, enabling novice developers to follow established best practices during quantum software development [57]. This SMS highlights that the literature on QCaaS reports five patterns, namely the API Gateway, Layered Architecture, Classic-Quantum Split, Service Façade, and Repository Pattern as shown in Figure 5. Figure 6 depicts pattern thumbnails [58] that provide a concise visual summary of the pattern, typically including its name and structural representation of pattern elements and their interactions.
@@
Example: API Gateway for Amazon Braket enables clients to access QC resources via a Gateway – a pattern used by Amazon as a platform agnostic solution to intermediate between QC users and QC vendors via Amazon Braket [59]. For example, in Figure 6 a), the thumbnail suggests a pattern named Quantum API Gateway as a layered system comprises of three main elements including a) Quantum Service Clients that can invoke available services on b) Quantum Server via routing and recommendation by c) Quantum API Gateway, detailed in [S1]. Figure 6 a) illustrates Amazon Braket that is a managed service provided by Amazon Web Services (AWS), using the API Gateway pattern, for accessing QC resources via QCaaS. The pattern comprises three key layers. At the client node, users aim to access QC resources, having Jupyter Notebook and AWS Console, enabling them to design, simulate, and submit quantum programs via the Braket SDK. The gateway service, i.e., Amazon Braket, acts as an intermediary that abstracts hardware complexity and manages execution workflows, including device selection, queuing, and error handling. Finally, the quantum computing resources layer connects users to multiple quantum hardware providers (e.g., Vendor A, Vendor B, etc.). This model provides a platform-independent interface, facilitating seamless experimentation across various quantum technologies. By decoupling development from hardware integration, Amazon Braket supports scalable, vendor-neutral QC for researchers, developers, and organisations exploring quantum applications.
@@
Pattern documentation: One of the most common mechanisms to document and store patterns are pattern templates and pattern catalogues that provide structured management of available patterns [58]. Our intent is not to provide comprehensive details on pattern documentation and management, therefore, we rely on pattern thumbnails from Figure 6 to concisely summarise the identified patterns. We briefly introduce each pattern and elaborate on the quantum-classic split pattern, with an example below.
@@
  • Quantum-Classic Split is identified in a total of 3 studies is a quantum version of the Splitter pattern and enables hybrid computation that can be divided between classical and quantum computations [S10][12][s24]. As shown in Figure 6 c), the classic-quantum split is driven by the workflow, also referred to as the split logic to divide the execution of computations. Figure 6 shows that the qunatum-classic workflow modeled as a UML activity diagram computes and delegates activities to be executed on classical and quantum computers. For example, the component and connector architectural view in Figure 6 e) shows that a ClassicalComponent and QuantumDriver deployed on a classical machine can invoke QuantumComponent on a quantum computer. A typical real-world example of the quantum-classic split pattern is Shor’s algorithm for prime factorisation. The algorithm divides tasks between classical and quantum computations such that classical processing handles choosing a random integer, and performing the greatest common divisor checks, and postprocessing, whereas the quantum part efficiently determines the period using modular exponentiation and Quantum Fourier Transform (QFT). This hybrid approach leverages quantum speedup for order finding while relying on classical hardware for the remaining computations.
@@
  • Service Façade for quantum software services refers to a classical interface that abstracts and simplifies access to complex quantum functionalities. It acts as a bridge between classical clients and quantum backends, hiding the complexities of quantum computing, such as circuit design, QuBit management, and backend selection, behind a well-organised and easy to implement service-oriented interfaces. Service façade is identified into a total of 2 studies. For example, the study [S7] utilises the service façade pattern, which wraps the classical (legacy software services) under the interfaces of quantum microservices that can be executed on Amazon Braket platform.
@@
L Quantum service modeling supports a model-based representation (i.e., creating design or architecture models) of functionality and quality of the conceived services. Modeling phase is supported via: – Modelling notations can assist software engineers to shift their focus from complex implementation (low-level source-coding) specific details to a design perspective (high-level system view). Modeling can incrementally translate functional aspects into service models, enabling service implementation through model-driven engineering or low-code development. The literature highlights that this modeling phase should utilise both existing and innovative notations, such as UML diagrams and quantum circuit models, to design quantum software services. – Patterns complementăsoftware models by providing reuse knowledge and best practices of classical and quantum SE during service modelling. The review has identified several classical as well as quantum-specific patterns such as Quantum API Gateway and Qunatum-Classic Split that can facilitate architects to design quantum software services.